OpenStudy (anonymous):

in 2005 , 8.1 x 10 to the 10th power text messages were sent in the united states.by 2007,the number of annual text messages had risen to 3.63 X 10 to the 11th power . about how many times as great was the number of text messages in 2007 than 2005?

OpenStudy (radar):

Looks like a division is needed 3.63 10^11 ---------- = 0.448 10^1 =4.48 or texting in 2007 was 4.48 times that of 2005 8.1 10^10

OpenStudy (anonymous):

thank u

OpenStudy (radar):

My line was too long, imagine that 2005 is the last line not scrunched in above the denominator lol. do I need to re do it.

OpenStudy (radar):

3.63 10^11 ---------- = 0.448 10^1 =4.48 or 8.1 10^10 texting in 2007 was 4.48 times that of 2005